OpenStudy (anonymous):

I am completely lost on the entire subject of trigonometric functions. I missed majority of class when we discussed it first, and I have no idea what I'm doing when it comes to inverses and domains... Help?

OpenStudy (anonymous):

...awaiting a question needing assistance.

OpenStudy (anonymous):

best to read about it in a text or ask a specific question here. here is a decent cheat sheet that might help

OpenStudy (anonymous):

Can you explain how to determine the period of a function?

OpenStudy (anonymous):

it is hard (impossible) to write a complete explanation here. try a problem, and then if you are stuck ask a specific question.

OpenStudy (anonymous):

Period = time of complete cycle

OpenStudy (anonymous):

if you have a graph you can see the period from the graph the period of sine and cosine is \(2\pi\) for tangent it is \(\pi\) and if you have \[\sin(bx)\] it is \(\frac{2\pi}{b}\)

OpenStudy (anonymous):

the function would need to be a repeating one, like a waveform

OpenStudy (anonymous):

Okay, so a period is similar to a wavelength?

OpenStudy (anonymous):

Not per se, Sine wave goes above zero, then below, then reaches zero again from start to finish, that is the period

OpenStudy (anonymous):

wavelength is sort of different

OpenStudy (anonymous):

Oh, okay, that's very helpful. Also, how do you determine what the domain will be?

OpenStudy (anonymous):

generally, a repeating wave is measured from +peak to +peak

OpenStudy (anonymous):

here is a graph of \(\sin(2x)\) you can see that the period is \(\pi\) http://www.wolframalpha.com/input/?i=sin%282x%29

OpenStudy (anonymous):

Well, anyway, thank you all. You're all very helpful :-)

OpenStudy (anonymous):

You're welcome!

OpenStudy (australopithecus):

Here is an old note from my calc highschool class that explains the parts of a trig function so you can graph it. Don't know if it helpful but yeah